Question two:

(a) Harcourt Ltd bought furniture on hire purchase from Nkalimeng Furnitures on January 2017. If Harcourt pays the full amount immediately, he will pay $120,000. Instead, he decided to make down payment of $25,000, and pay the remaining balance in instalments. The first instalment of $12,000 was paid in December 2017. The next 3 instalments of $20,500 were also paid at the end of the 2018, 2019 and 2020 respectively. He had to pay the balance in 2021.

Required: calculate the total interest paid by Harcourt Ltd.


(b) Create a scenario to demonstrate how Hire Vendor will account for the items bought under

hire purchase. Hint: your scenario must include the following:

i)Asset a/c

ii)Suppliers a/c

iii)Hire purchase interest a/c

iv)Statement of financial position extract

v)The instalments should be paid three times


1
Expert's answer
2022-02-16T15:03:01-0500

a)

let the rate be 3%

120 000-25 000=95000

n=5

"A=95 000\times(0.03+\frac{0.03}{(1+0.03)^{60}-1})=3432.63"

"3432.63\times60=205957.9"

the total interest paid:

205957.9-95000=110957,9


b)On the day of shipment, the seller's entry in the debit of the account "Settlements with buyers and customers" and the credit of the account "Sales" reflects revenue in the amount of the contractual value of the goods. Interest accrued on the last day of each following month (starting from the month of the loan) can increase revenue (a similar posting is made). Or you can account for them as other income, which is more convenient, because with this approach, accounting will be closer to tax accounting: debit account Settlements with buyers and customers - credit account Other income and expenses ”. Any of the proposed options must be registered in the accounting policy of the organization.

Upon receipt of the goods, the buyer only takes into account its value under the contract. And interest will be reflected in accounting as it accrues on the last day of each month and on the date of repayment of the loan as part of other expenses (debit of the account “Other income and expenses”, - credit of the account “Settlements with suppliers and contractors”